A chemistry student mixes together 12mL of a 10% chlorine solution, 5 mL of 80% chlorine solution and 20mL of a 45% solution. About what percent chlorine is the final solution?
Answer by mananth(12270)

12mL of a 10% ,
5 mL of 80%
20mL of a 45%
..
total quantity of mixture = 37 ml of x%
..
concentration equation
0.1*12 +0.8*5+0.45*20 = 0.08x*37
1.2+4+9= 2.96x
14.2= 2.96x
x= 14.2/2.96
x= 4.8

% = 38